# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v3

@@ 3.2 Effective Connectivity Model
$\tau_{x\rightarrow y}(t)=\operatorname*{arg\,max}_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$ (9)
Hence, if $\tau_{x\rightarrow y}(t)=0$, there is a zero-propagation delay for the information flow going from $x(t)$ to $y(t)$.
3.3 Desynchronization Index
To identify ictal channels from the connectivity analysis, we follow a similar approach to that presented in [42], where Sparks and colleagues extended the Exponentially Weighted Moving Average (EWMA) statistic to detect anomalous behaviors in time-varying networks. In particular, we assume that the generation of epileptic discharges is associated with an abrupt reduction in the communication between the EZ and the rest of the cortical structures.
@@ 3.3 Desynchronization Index
$\mathcal{E}(\omega,t)=\begin{cases}\alpha\omega(t)+(1-\alpha)\mathcal{E}(\omega,t-\Delta t)\;&t>0;\\ \omega(t)\;&t=0;\end{cases}$ (10)
@@
[…]
In other words, we consider abnormal those connections that go outside the inter-quartile range of the time-varying network distribution. This, on the one hand, allows us to implicitly customize the framework to each patient without the need to define additional parameters as in [42]. On the other hand, choosing other percentile ranges for discerning abnormal connections could lead to different trade-offs in terms of accuracy, and, in the future, further investigation of how to tune such a parameter is required.
Table 1: Demographic and clinical data.
Patient index Sex Age (SEEG) Age (onset) Seizure frequency EZ localization PZ localization Surgery Outcome (1 year) Histology
1 Male 39 33 Weekly Right temporal Temporal pole Lobectomy Seizure free HS 1
@@ 2 Male 24 1 Daily Left orbito-frontal Fronto-insular RF-TC Seizure reduction Not available
3 Male 36 16 Monthly Left temporal Temporal pole Lobectomy Seizure free FCD 1b
@@ 4 Female 52 3 Weekly Right temporal Temporal and FOI Lobectomy Seizure free HS 1
5 Female 25 2 Weekly Right temporo-basal Temporo-parietal and insular Lobectomy Seizure free Aspecific
6 Male 48 22 Weekly Right temporal Orbito-frontal, and insular Lobectomy Seizure free HS 1
7 Male 36 9 Weekly Right temporo-occipital Temporo-basal and mesial RF-TC Seizure reduction Not available
@@ 8 Female 49 12 Monthly Left temporal Temporo-lateral Lobectomy Seizure free HS 1
9 Male 24 18 Weekly Right temporal Temporal pole RF-TC Seizure free Tumor
@@ 10 Female 38 13 Monthly Right temporo-occipital Temporo-mesial RF-TC Not available Not available
11 Male 32 1 Monthly Left temporo-occipital Parietal and insulo-opercular RF-TC Not available Not available
@@
$\psi_{x}^{\text{in}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{in}}(t)}\mathcal{E}(T_{y\rightarrow x},t),$ (12)
$\hat{\psi}_{x}^{\text{in}}(t)=|\mathcal{N}_{x}^{\text{in}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$ (13)
In a specular way, we define $\psi_{x}^{\text{out}}(t)$ and $\hat{\psi}_{x}^{\text{out}}(t)$ as
$\psi_{x}^{\text{out}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{out}}(t)}\mathcal{E}(T_{x\rightarrow y},t),$ (14)
$\hat{\psi}_{x}^{\text{out}}(t)=|\mathcal{N}_{x}^{\text{out}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$ (15)
We now define the desynchronization level of channel $x\in\mathcal{N}$ at time $t$ as
$D_{x}(t)=\sqrt{\hat{\psi}_{x}^{\text{in}}(t)}-\sqrt{\psi_{x}^{\text{in}}(t)}+\sqrt{\hat{\psi}_{x}^{\text{out}}(t)}-\sqrt{\psi_{x}^{\text{out}}(t)},$ (16)
where the square root operator is used to stabilize the variance of the process [42]. The first part of the above equation depends on the inward connection of $x$, while the second refers to the influence that $x$ exerts on other channels. At each time $t$, $D_{x}(t)$ represents the tendency of $x$ to self-isolate from the rest of the network.
@@
$DI_{x}=\begin{cases}\frac{c_{D_{x}}}{t_{D_{x}}-t_{\text{start}}}&x\in\mathcal{N}_{D},\\ 0&x\notin\mathcal{N}_{D}.\end{cases}$ (17)
@@ 4.1 Clinical Dataset
The SEEG implantation followed the workflow developed at Niguarda Hospital [43], which involves the construction of a multimodal scene of the patient’s brain. The scene allows for a comprehensive evaluation of all the anatomical information regarding the cortical area explored by each contact. The SEEG signals were recorded using the Nihon Kohden 2100 polygraph, considering 192 or 256 channels and a sampling frequency of $f_{s}=1000$ Hz. High-definition synchronized videos were recorded for the whole duration of each SEEG monitoring (up to 20 days per patient), enabling a correlation between electrical and clinical features. Finally, we collected the main demographic and clinical data of each patient, including age at disease onset, seizure semiology, neuropsychological examinations, neuroimaging findings, scalp EEG recordings, surgical outcomes, follow-up information, and histopathological diagnoses.
From the initial dataset, a single patient was excluded as the SEEG record did not allow the localization of the EZ and no epilepsy surgery was performed. The final cohort comprises 7 males and 4 females with an average age of 36.6 years (range 24-49) at the recording time, and 8.8 years (range 1-22) at the disease onset. Clinical data indicated a probable EZ in the temporal lobe in 7 cases (5 right, 2 left), in the left frontal lobe in a single case, and in the temporo-occipital lobe in the remaining cases (2 right, 1 left). Radiofrequency thermocoagulations allowed control of seizures in a single patient, while the remaining patients underwent anterior-temporal lobectomy. Histopathological analysis led to clinical results compatible with hippocampal sclerosis in 4, and cortical dysplasia in 1 cases. Demographic and clinical details of the studied population are reported in Tab. 1  [^1] HS stays for hippocampal sclerosis FCD for focal cortical dysplasia, RF-TC for radio-frequency thermocoagulation, FOI for fronto-orbito-insular..
4.2 Performance Evaluation
@@
Before running the EI and DI algorithms, the SEEG signals were pre-processed through a notch filter centered at $f_{\text{notch}}=50$ Hz. In the analysis, we excluded sites that explore the White Matter (WM) since they are not sources of electric signals. The selection of the WM channels was performed by a board-certified neurophysiologist (L.F.), according to the electrical activity of the cortical sites and their location on the multimodal scene. No other cortical sites were excluded, making our framework agnostic to the specific patient and seizure to be analyzed.
@@
In addition, we study the trade-off between sensitivity and precision by computing the Receiver Operating Characteristic (ROC) and the Area Under the ROC Curve (AUC). The AUC is a common performance indicator for diagnostic tools: in clinical settings, a value higher than $0.80$ is solely required for obtain a reliable solution [44]. Practically, for each patient, the ROC curve is obtained by varying the number $M$ of epileptogenic channels up to the size of the SEEG implant.
The EI and DI algorithms are both implemented as standalone tools or combined according to the “and” and “or” Boolean functions. Practically, in the first case (named “EI and DI”), a channel is classified as epileptogenic only if both the algorithms mark it as positive. Instead, in the second case (named “EI or DI”), a channel is classified as epileptogenic if at least one of the algorithms classifies it as positive. Clearly, we expect the first configuration to provide higher sensitivity, and the second configuration to return less false positives.
@@ 5 Results
5.1 Detection Results
The SEEG signals are segmented in time windows lasting $T_{\text{window}}=1.00$ s each, considering a time shift of $\Delta t=0.25$ s. Given the epoch duration ($T_{\text{epoch}}=200$ s) a total of $797$ windows is analyzed for each seizure. For estimating the connectivity between different channels according to the PTE algorithm, we consider $\tau_{max}=0.10$ s as maximum propagation delay. Instead, for computing the energy ratio, the high and low frequency ranges are set to $B_{h}=[30.0,250.0]$ Hz and $B_{\ell}=[4.0,12.0]$ Hz, respectively. Such a setting extends the portion of the spectrum processed by the EI algorithm, which, in its initial version, considered frequencies lower than $97$ Hz [8]. Here, we consider a wider frequency range to capture ripple phenomena that are notably characterized by oscillations up to $250$ Hz or even more [45].
We set the tonicity interval to $\delta=5.00$ s (as done in [8]), the weight of the CUSUM chart to $\gamma=0.0$, and the exponential decay of the EWMA chart to $\alpha=1.0$, following the recommendations provided in [46]. We set the detection threshold of the algorithms to $\eta=0.0$, which means that no channels were discarded from the algorithm output and the number of epileptogenic channels is always equal to $M$. Finally, we set the baseline period of the CUSUM chart to $t_{\text{base}}=t_{\text{start}}-20.0$ s so that the energy and connectivity statistics of each channel are estimating according to the $20.0$ s preceding the onset of the seizure. As explained in Sec. 4, $t_{\text{start}}$ and $t_{\text{end}}$ depend on the specific seizure analyzed and, therefore, are epoch-tailored.
@@ 5.1 Detection Results
Figure 1: Performance metrics in the EZ detection.
Figure 2: Area Under the ROC Curve (AUC) in the EZ detection.
In Fig. 1(a), we report the detection performance of the different algorithms, considering the channels within the EZ as a classification target. In particular, we tune $M$ is such a way $10\%$ of the channels are considered epileptogenic. We can appreciate that DI has higher sensitivity than EI ($0.44$ and $0.40$, respectively), while the ”EI or DI” configuration increases the sensitivity more than $50$% with respect EI (with a final score of $0.60$). This denotes that two algorithms return different results, i.e., the channels associated with a higher EI value are not the same associated with a higher DI value. Focusing on precision, the DI algorithm still overcomes EI ($0.40$ vs $0.34$), while considering the “and” function leads to a performance gain of $20\%$. Overall, the best accuracy is obtained using the “EI or DI” approach, which returns a score of $0.95$, followed by the DI algorithm at second place.
To improve sensitivity, we can increase the number $M$ of channels returned by the algorithms. When we set $M=20\%$, the sensitivity of the “EI or DI” approach reaches a value of $0.77$ while the accuracy decreases to $0.94$ due to the higher number of false positives (Fig. 1(b)). On the other hand, we highlight that the main advantage of our technique is to build first sketch of EZ, allowing neurophysiologists to look at a limited number of channels relative to the original network. This is strongly beneficial in the case of extensive SEEG implants (with hundreds of channels), where understanding the time-varying dynamics that lead to seizures is extremely difficult with a pure visual analysis.
Fig. 2 denotes how the ”EI or DI” approach leads to the best detection performance ($AUC=0.87$), followed by the DI algorithm ($AUC=0.80$) and the EI algorithm ($AUC=0.74$) at third place. These results confirm that the combined use of EI and DI enables a more holistic analysis of the overall ictal transition, allowing us to identify most of the channels that are part of the EZ. On the other hand, to ensure a sensitivity close to $0.8$, we must still look at more than $M=20\%$ of the SEEG channels (as we observed in Fig. 1(b)). We expect that integration of the proposed tools with other quantitative features, for example, taking into account the space position of recording sites within the implant, can further enhance the detection performance. In a clinical setting, this will make it possible to further reduce the neurophysiologist effort, i.e., the number $M$ of channels to be individually evaluated.
@@
Figure 3: SEEG implants.
5.2 Concordant Example
In the following, we better evaluate the operations of the two algorithms, analyzing two specific seizure events. We first consider patient $7$, whose SEEG implant (represented in Fig. 3(a)) explored the right temporo-parieto-occipital area, and includes $152$ different recording sites ($18$ electrodes). In Fig. 4(a), we report the bipolar representation of the $M=10$ channels associated with the strongest variation in the energy ratio during the 50 seconds preceding the propagation of epileptic discharges. Instead, in Fig. 4(b), we report the energy ratio values associated with each channel.

### v4

@@ 3.2. Effective Connectivity Model
(9) $\tau_{x\rightarrow y}(t)=\operatorname*{arg\,max}_{\tau\in[0,\tau_{max}]}T_{x\rightarrow y}(t,\tau).$
Hence, if $\tau_{x\rightarrow y}(t)=0$, there is a zero propagation delay for the information flow going from $x(t)$ to $y(t)$.
3.3. Desynchronization Index
To identify ictal channels from the connectivity analysis, we follow a similar approach to that presented in (Sparks and Wilson, 2019), where Sparks and colleagues extended the Exponentially Weighted Moving Average (EWMA) statistic to detect anomalous behaviors in time-varying networks. In particular, we assume that the generation of epileptic discharges is associated with an abrupt reduction in the communication between the EZ and the rest of the cortical structures.
@@ 3.3. Desynchronization Index
(10) $\mathcal{E}(\omega,t)=\begin{cases}\alpha\omega(t)+(1-\alpha)\mathcal{E}(\omega,t-\Delta t)\;&t>0;\\ \omega(t)\;&t=0;\end{cases}$
@@
(11) $\mathcal{N}_{x}^{\text{in}}(t)$ =$=\big{\{}y\in\mathcal{N}:\mathcal{E}(T_{y\rightarrow x},t)\leq\mathcal{E}(P_{25}[\mathcal{T}],t)\big{\}},$
$\mathcal{N}_{x}^{\text{out}}(t)$ =$=\big{\{}y\in\mathcal{N}:\mathcal{E}(T_{x\rightarrow y},t)\geq\mathcal{E}(P_{75}[\mathcal{T}],t)\big{\}}.$
In other words, we consider abnormal those connections that go outside the inter-quartile range of the time-varying network distribution. This, on the one hand, allows us to implicitly customize the framework to each patient without the need to define additional parameters as in (Sparks and Wilson, 2019). On the other hand, choosing other percentile ranges for discerning abnormal connections could lead to different trade-offs in terms of accuracy, and, in the future, further investigation of how to tune such a parameter is required.
Table 1. Demographic and clinical data.
Patient index Sex Age (SEEG) Age (onset) Seizure frequency Epileptogenic zone localization Propagation zone localization Surgery Outcome (1 year) Histology
1 Male 39 33 Weekly Right temporal Temporal Lobectomy Seizure free HS 1
@@ 2 Male 24 1 Daily Left orbito-frontal Fronto-insular RF-TC Seizure reduction Not available
3 Male 36 16 Monthly Left temporal Temporal Lobectomy Seizure free FCD 1b
@@ 4 Female 52 3 Weekly Right temporal Temporal and FOI Lobectomy Seizure free HS 1
5 Female 25 2 Weekly Right temporo-basal Parietal and insular Lobectomy Seizure free Aspecific
6 Male 48 22 Weekly Right temporal Fronto-orbito-insular Lobectomy Seizure free HS 1
7 Male 36 9 Weekly Right temporo-occipital Temporal RF-TC Seizure reduction Not available
@@ 8 Female 49 12 Monthly Left temporal Temporo-lateral Lobectomy Seizure free HS 1
9 Male 24 18 Weekly Right temporal Temporal RF-TC Seizure free Tumor
@@ 10 Female 38 13 Monthly Right temporo-occipital Temporo-mesial RF-TC Not available Not available
11 Male 32 1 Monthly Left temporo-occipital Insulo-opercular RF-TC Not available Not available
@@
(12) $\psi_{x}^{\text{in}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{in}}(t)}\mathcal{E}(T_{y\rightarrow x},t),$
(13) $\hat{\psi}_{x}^{\text{in}}(t)=|\mathcal{N}_{x}^{\text{in}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$
In a similar way, we define $\psi_{x}^{\text{out}}(t)$ and $\hat{\psi}_{x}^{\text{out}}(t)$ as
(14) $\psi_{x}^{\text{out}}(t)=\sum_{y\in\mathcal{N}_{x}^{\text{out}}(t)}\mathcal{E}(T_{x\rightarrow y},t),$
(15) $\hat{\psi}_{x}^{\text{out}}(t)=|\mathcal{N}_{x}^{\text{out}}(t)|\cdot\mathcal{E}(M[\mathcal{T}],t).$
Finally, we define the desynchronization level of channel $x\in\mathcal{N}$ at time $t$ as
(16) $D_{x}(t)=\frac{\sqrt{\hat{\psi}_{x}^{\text{in}}(t)}-\sqrt{\psi_{x}^{\text{in}}(t)}}{\sqrt{\psi_{x}^{\text{out}}(t)}-\sqrt{\hat{\psi}_{x}^{\text{out}}(t)}},$
where the square root operator is used to stabilize the variance of the process (Sparks and Wilson, 2019). The numerator of the above equation depends on the inward connection of $x$, while the denominator refers to the influence that $x$ exerts on other channels. Hence, $D_{x}(t)$ represents the tendency of $x$ to self-isolate from the rest of the network.
@@
(17) $DI_{x}=\begin{cases}\frac{c_{D_{x}}}{t_{D_{x}}-t_{\text{start}}}&x\in\mathcal{N}_{D},\\ 0&x\notin\mathcal{N}_{D}.\end{cases}$
@@ 4.1. Clinical Dataset
The SEEG implantation followed the workflow developed at Niguarda Hospital (Cardinale et al., 2015), which involves the construction of a multimodal scene of the patient’s brain. The scene allows for a comprehensive evaluation of all the anatomical information regarding the cortical area explored by each contact. The SEEG signals were recorded using the Nihon Kohden 2100 polygraph, considering 192 or 256 channels and a sampling frequency of $f_{s}=1000$ Hz. High-definition synchronized videos were recorded for the whole duration of each SEEG monitoring (up to 20 days per patient), enabling a correlation between electrical and clinical features. Finally, we collected the main demographic and clinical data of each patient, including age at disease onset, seizure semiology, neuropsychological examinations, neuroimaging findings, scalp EEG recordings, surgical outcomes, follow-up information, and histopathological diagnoses.
From the initial dataset, a single patient was excluded as the SEEG record did not allow the localization of the EZ and no epilepsy surgery was performed. The final cohort comprises 7 males and 4 females with an average age of 36.6 years (range 24-49) at the recording time, and 8.8 years (range 1-22) at the disease onset. Clinical data indicated a probable EZ in the temporal lobe in 7 cases (5 right, 2 left), in the left frontal lobe in a single case, and in the temporo-occipital lobe in the remaining cases (2 right, 1 left). Radiofrequency thermocoagulations allowed control of seizures in a single patient, while the remaining patients underwent anterior-temporal lobectomy. Histopathological analysis led to clinical results compatible with hippocampal sclerosis in 4, and cortical dysplasia in 1 cases. Demographic and clinical details of the studied population are reported in Tab. 1  [^1] In the table, HS stays for hippocampal sclerosis FCD for focal cortical dysplasia, RF-TC for radio-frequency thermocoagulation, FOI for fronto-orbito-insular..
4.2. Performance Evaluation
@@
Before running the EI and DI algorithms, the SEEG signals were pre-processed through a comb filter centered at $f_{\text{comb}}=50$ Hz to remove the powerline frequency and its harmonics. In the analysis, we excluded sites that explore the White Matter (WM) since they are not sources of electric signals. The selection of the WM channels was performed by a board-certified neurophysiologist (L.F.), according to the electrical activity of the cortical sites and their location on the multimodal scene. No other cortical sites were excluded, making our framework agnostic to the specific patient and seizure to be analyzed.
@@
In addition, we study the trade-off between sensitivity and precision by computing the Receiver Operating Characteristic (ROC) and the Area Under the ROC Curve (AUC). The AUC is a common performance indicator for diagnostic tools: in clinical settings, a value higher than $0.80$ is solely required for obtain a reliable solution (Hosmer Jr et al., 2013). Practically, for each patient, the ROC curve is obtained by varying the number $M$ of epileptogenic channels up to the size of the SEEG implant.
The EI and DI algorithms are both implemented as standalone tools or combined according to the “and” and “or” Boolean functions. Practically, in the first case (named “EI and DI”), a channel is classified as epileptogenic only if both the algorithms mark it as positive. Instead, in the second case (named “EI or DI”), a channel is classified as epileptogenic if at least one of the algorithms classifies it as positive. Clearly, we expect the first configuration to return less false positives, and the second configuration to provide higher sensitivity.
@@ 5. Results
5.1. Detection Results
The SEEG signals are segmented in overlapping time windows lasting $T_{\text{window}}=1.00$ s each, considering a time shift of $\Delta t=0.25$ s between consecutive windows. Given the epoch duration ($T_{\text{epoch}}=200$ s) a total of $797$ windows is analyzed for each seizure. To estimate the connectivity between different channels according to the PTE algorithm, we consider $\tau_{max}=0.10$ s as the maximum propagation delay. Instead, for computing the energy ratio, the high and low frequency ranges are set to $B_{h}=[30.0,250.0]$ Hz and $B_{\ell}=[4.0,12.0]$ Hz, respectively. This setting extends the portion of the spectrum processed by the EI algorithm, which, in its initial version, considered frequencies lower than $97$ Hz (Bartolomei et al., 2008). Here, we consider a wider frequency range to capture ripple phenomena that are notably characterized by oscillations up to $250$ Hz or even more (Urrestarazu et al., 2007).
We set the tonicity interval to $\delta=5.00$ s (as done in (Bartolomei et al., 2008)), the weight of the CUSUM chart to $\gamma=0.0$, and the exponential decay of the EWMA chart to $\alpha=1.0$, following the recommendations provided in (Montgomery, 2019). We set the detection threshold of the algorithms to $\eta=0.0$, which means that no channels were discarded from the algorithm output and the number of epileptogenic channels is always equal to $M$. Finally, we set the baseline period of the CUSUM chart to $t_{\text{base}}=t_{\text{start}}-20.0$ s so that the energy and connectivity statistics of each channel are estimating according to the $20.0$ s preceding the onset of the seizure. As explained in Sec. 4, $t_{\text{start}}$ and $t_{\text{end}}$ depend on the specific seizure analyzed and, therefore, are epoch-tailored.
@@ 5.1. Detection Results
Figure 1. Performance metrics in the EZ detection.
[…]
[Performance metrics in the EZ detection.]Performance metrics in the EZ detection.
Figure 2. Area Under the ROC Curve (AUC) in the EZ detection.
[…]
[Area Under the ROC Curve (AUC) in the EZ detection.]Area Under the ROC Curve (AUC) in the EZ detection.
@@
Figure 3. SEEG implants.
[…]
[SEEG implants.]SEEG implants.
In Fig. 1(a), we report the detection performance of the different algorithms, considering the channels within the EZ as a classification target. In particular, we tune $M$ is such a way $10\%$ of the channels are considered epileptogenic. We note that DI has higher sensitivity than EI ($0.56$ and $0.35$, respectively), while the ”EI or DI” configuration increases the sensitivity more than $50$% with respect EI (with a final score of $0.65$). This denotes that two algorithms return different results, i.e., the channels associated with a higher EI value are not the same associated with a higher DI value. Focusing on precision, the DI algorithm still overcomes EI ($0.45$ vs $0.33$), while considering the “and” function does not improve the performance than when using DI. The best accuracy is obtained using the “EI or DI” approach, which returns a score of $0.95$, followed by the DI algorithm at second place.
To improve sensitivity, we can increase the number $M$ of channels returned by the algorithms. When we set $M=20\%$, the sensitivity of the “EI or DI” approach reaches a value of $0.79$ while the accuracy decreases to $0.92$ due to the higher number of false positives (Fig. 1(b)). On the other hand, we highlight that the main advantage of our technique is to build a first sketch of EZ, allowing neurophysiologists to look at a limited number of channels relative to the original network. This may be strongly beneficial in the case of extensive SEEG implants (with hundreds of channels), where understanding the time-varying dynamics that lead to seizures is extremely difficult with pure visual analysis.
Fig. 2 denotes how the ”EI or DI” approach leads to the best detection performance ($AUC=0.87$), followed by the DI algorithm ($AUC=0.81$) and the EI algorithm ($AUC=0.74$) at third place. These results confirm that the combined use of EI and DI enables a more holistic analysis of the overall ictal transition, allowing us to identify most of the channels that are part of the EZ. On the other hand, to ensure a sensitivity close to $0.8$, we must still look at more than $M=20\%$ of the SEEG channels (as we observed in Fig. 1(b)). We expect that integration of the proposed tools with other quantitative features, for example, taking into account the space position of recording sites within the implant, can further enhance detection performance. In a clinical setting, this will make it possible to further reduce the neurophysiologist effort, i.e., the number $M$ of channels to be individually evaluated.
5.2. Concordant Example